# RAUCA: A Novel Physical Adversarial Attack on Vehicle Detectors via Robust and Accurate Camouflage Generation

### v1

@@ 1 Introduction
Figure 1: Comparison of different adversarial camouflage under daytime (first row) and night (second row) environments, where only our method succeeds in both cases. (a) A car with normal texture; (b) DAS (Wang et al., 2021). (c) and (d) are top-performed methods FCA (Wang et al., 2022) and ACTIVE (Suryanto et al., 2023), respectively. (e) Our method RAUCA.
@@
Deep Neural Networks (DNNs) have achieved remarkable performance in many real-world applications such as face recognition and autonomous vehicles (Krizhevsky et al., 2012; Wang et al., 2023; Chai et al., 2021). However, DNNs suffer from adversarial example attacks (Szegedy et al., 2014). For instance, in the task of vehicle detection, adversarial inputs can deceive detection models, leading to incorrect detection of the surrounding vehicles, which poses a severe threat to the safety of autonomous vehicles.
@@
Adversarial example attacks can be categorized into digital and physical attacks. Digital attacks involve introducing pixel-level perturbations to the model’s digital input, while physical attacks manipulate real-world objects or surroundings, indirectly influencing the model inputs (Chen et al., 2018; Sharif et al., 2016a; Mahendran & Vedaldi, 2015). Physical attacks are generally deemed more practical, as gaining direct access to the model’s input often necessitates system authentication. However, such attacks are inherently more challenging as they must prove effective in complex physical environments, including various viewing perspectives, spatial distances, and lighting/weather conditions.
@@
This paper focuses on physical attacks against vehicle (e.g., car) detection models, given their wide adoption in autonomous driving scenarios where safety is of great importance. To ensure the attack effectiveness across various viewing angles, current methods prefer generating adversarial camouflage capable of covering the entire surface of the vehicle (Wang et al., 2021, 2022; Suryanto et al., 2022, 2023). Top-performing methods achieve this by leveraging a differential neural renderer. This renderer maps the 3D vehicle and its texture to 2D images, establishing a differentiable path between the 3D vehicle and the vehicle detection models. Thus, the texture can be optimized through gradient back-propagation for effective camouflage generation.
@@
In the literature, there are two ways to generate camouflage with a neural renderer: one way is to optimize a 2D square texture pattern and project it onto the vehicle repeatedly, referred to as world-align-based methods; the other is to optimize the 3D texture of the vehicle in the form of UV-maps, referred to as UV-map-based method. However, both approaches currently suffer from certain issues. The former method cannot project the texture pattern to the car during evaluation to match the way used in optimization precisely, leading to differences in adversarial camouflage between generation and evaluation. The latter UV-map-based render methods are not capable of rendering sophisticated environment characteristics such as light and weather, leading to sub-optimal camouflage generation. Besides, all current methods fail to consider the effectiveness of the camouflage under various weather conditions. As shown in Figure 1, the camouflage generated by the top-performed FCA (Wang et al., 2022) and ACTIVE (Suryanto et al., 2023) methods fails to attack the detector at night.
@@
Our extensive studies demonstrate that our method outperforms the current state-of-the-art method by around 10% in car detection performance (e.g., AP@0.5). Our attacks achieve strong robustness and effectiveness, excelling in multi-weather, multi-view, and multi-distance conditions. We also demonstrate our method is effective in various weather conditions in the physical world. Our code is available at: https://anonymous.4open.science/r/RAUCA-6F34.
@@ 2 Related work
Adversarial Camouflage: For self-driving cars, precise detection of surrounding vehicles is a critical safety requirement. Consequently, there has been a growing interest in developing adversarial vehicle camouflage to evade vehicle detection systems. Current research works mostly leverage a 3D simulation environment to obtain 2D rendered vehicle images with various transformations such that they can obtain robust adversarial camouflage. The early researches of adversarial camouflage for vehicles are mostly black-box methods because the rendering process in the simulation environment is non-differentiable. Zhang et al. (2019) first conducts experiments in the 3D space to generate adversarial camouflage for cars. They propose CAMOU, a method to train an approximate gradient network to mimic the behavior of both rendering and detection of the camouflage vehicles. Then, they can optimize the adversarial texture using this network. Meanwhile, Wu et al. (2020) propose an adversarial camouflage generation framework based on a genetic algorithm to search an adversarial texture pattern. Then they repeat the optimized texture pattern to build the 3D texture that covers the full-body vehicle.
Recent methods introduce neural renderers, which enable differentiable rendering. With this technique, the adversarial texture can be optimized via gradient back-propagation. Two primary methods are employed in adversarial camouflage generation with neural renderer. One is to optimize a 3D model texture directly, which we refer to as UV-map-based camouflage methods; for example, Wang et al. (2021) propose a Dual Attention Suppression (DAS) attack, which minimizes the model attention and human attention on the camouflaged vehicle. Besides, Wang et al. (2022) propose the Full-coverage Camouflage Attack (FCA), which optimizes full-body surfaces of the vehicle in multi-view scenarios. The other is to optimize a 2D square texture pattern and then project it repeatedly to the target vehicle surface, which we refer to as world-align-based camouflage methods. Suryanto et al. (2022) present the Differentiable Transformer Attack (DTA), which proposes a differentiable renderer that can express physical and realistic characteristics (e.g. shadow). ACTIVE proposed by Suryanto et al. (2023) introduces a texture mapping technique that utilizes depth images and improves the naturalness of the camouflage by using larger texture resolution and background colors.
@@ 3.1 Overview
Figure 2 shows our entire framework for adversarial camouflage generation. Firstly, we modify the weather parameters in the simulation environment to obtain a multi-weather vehicle dataset ($I_{in}$, $Y$, $\Phi_{cam}$, $M$), where $I_{in}$ is the original input images, $Y$ is the ground truth labels, $\Phi_{cam}$ is the camera pose parameters (position and angle) for viewing the car, and $M$ is the binary masks of $I_{in}$, where the target vehicle areas are set to 1. With $I_{in}$ and $M$, we can use:
$B=I_{in}\cdot\left(1-{M}\right)$ (1)
$X_{ref}=I_{in}\cdot{M}$ (2)
@@ 3.3 Neural Render Plus (NRP)
Before applying the framework to generate adversarial textures, NRP needs to be trained in advance to optimize the parameters of EFE. The NRP training process inputs the masked images of the white car $X_{ref}$, 3D mesh $Msh$, poses of the camera $\Phi_{cam}$, and multiple preset 3D texture T with different colors. At the same time, we need to obtain images of different vehicle colors based on $\Phi_{cam}$ and the colors of T from the photo-realistic renderer. Then we cut out the vehicle parts as the network output target $TG$.
@@ 3.4 Attack Loss
where $x$ is the input image of the target detector, $H_{b}(x)$ is the detection bounding box, $gt$ is the ground-truth box and $IoU\left(H_{b}(x),gt\right)$ is the Intersection over Union (IoU) between $H_{b}(x)$ and $gt$. We use $IoU\left(H_{b}(x),gt\right)$ as a weight term, which allows the optimized loss function to focus more on the bounding box with a larger intersecting ratio area with $gt$. $H_{o}(x)$ and $H_{c}(x)$ represent the objectiveness score and the class confidence score for the bounding box, respectively. $H_{d}\left.(x\right.)$ is our detection score, which is the product of the objectiveness confidence, class confidence, and the intersect ratio. Due to $IoU\left(H_{b}(x),gt\right)$, we only assign a non-zero detection score to the detected boxes which are intersected with the ground-truth box. This makes the texture optimization focus on making the target vehicle detection ineffective. We select the highest $H_{d}(x)$ and use it to compute $L_{atk}(x)$ through a log loss. By minimizing $L_{atk}(x)$, we can make the camouflaged vehicle misclassified or undetected by the object detector.
@@ 4.1 Experimental Settings
Datasets: We utilize CARLA to generate datasets in our experimentation. To facilitate a comparative analysis with previous studies (Wang et al., 2021, 2022; Suryanto et al., 2022, 2023), we select the Audi E-Tron as the target vehicle model. We create datasets using various simulation settings. For NRP training and testing, we utilized 69,120 and 59,152 photo-realistic images, encompassing 16 distinct weather conditions, respectively. These conditions are a combination of four sun altitudes and four fog densities. Additionally, we generate a set of 40,960 images for our texture generation and prepare 8192 images for each adversarial camouflage in robust evaluation. The weather conditions present in these two datasets are the same conditions employed during the training and testing of NRP. For transferability evaluation, we utilize an “unseen-weather dataset” consisting of 8192 images in 16 novel weather conditions that are not seen during the texture generation phase. Moreover, we conduct experiments in the real world by printing five types of adversarial camouflages and sticking them to the 1:12 Audi E-Tron car models. We shoot 432 photos under different light conditions and 120 photos in a foggy environment for each model. Additional information regarding the construction process of the datasets can be found in the appendix.
@@ 4.2 Evaluation in Physically-Based Simulation Settings
Table 1: Comparison of the effectiveness of camouflages across various object detection models. Values are AP@0.5 of the car.
@@
Figure 4: Attack comparison in the foggy day, a weather condition that has been included in our training set.
Figure 4: Attack comparison in the foggy day, a weather condition that has been included in our training set.
@@
Robustness in multi-weather dataset: We run an extensive attack comparison using diverse detection models. In addition to YOLOv3, we also use various black-box detection models to evaluate the camouflage. The results are shown in Table 1, showing that RAUCA has the best performance on all models. DAS generally performs less effectively, primarily due to the limitations of partially painted camouflage. Meanwhile, FCA exhibits sub-optimal performance, only slightly better than random camouflage, because it is not capable of rendering sophisticated environment characteristics during texture generation. DTA and ACTIVE are also less effective than our method because they do not consider multi-weather conditions during texture generation. Besides, their texture projection is not consistent between generation and testing. Our method demonstrates a nearly 13% improvement in attack effectiveness compared to other models on the white-box YOLOv3 model. Additionally, it achieves an improvement of around 8% on most other black-box models. Figure 5 shows an example of different vehicle camouflage in a foggy scene: our method is still effective, while other methods fail to attack.
Figure 6 shows the summarized performance of each camera pose and weather parameter; values are car AP@0.5 averaged from the detectors used in Table 1. We can see that the texture produced by our method consistently outperforms other approaches in all camera transformations and weather conditions. Our method improves multi-view robustness over previous methods in every viewpoint, thanks to the UV mapping-based projection for our textures and effective rendering of the environmental characteristics in the foreground. In addition, due to the incorporating multi-weather dataset during camouflage generation, our method also has the best robustness to different weather environments.
@@
Transferability in unseen-weather dataset: We further evaluate the transferability of our camouflage under unseen weather conditions. As shown in Table 2, our method achieves the best attack performance on all types of detectors. Our method outperforms the previous state-of-the-art method, ACTIVE by 23% on the white-box model YOLOv3, and approximately 11% on other black-box models. Our camouflage shows better attack performance under new weather conditions than under the weather conditions used for texture generation. This highlights the transferability of our approach across diverse weather conditions. Figure 5 shows an example of different camouflage on rainy days. Notice that rainy weather is excluded from texture generation. Even though the rain blurs part of the texture, RAUCA uniquely succeeds in attacking the target detector.
@@ 4.3 Evaluation in Real-World Settings
We average the detection results for all environmental cases and the results are in the last column of Table 3, which shows that our method has the best attack results. It illustrates the robustness of our method in different physical environments.
@@ 4.4 Ablation Studies
Effectiveness of the multi-weather dataset and NRP for texture generation: We evaluate the effectiveness of incorporating multi-weather conditions into our dataset and our proposed rendering component, NRP, for texture generation. As can be seen in Table 5, we generate the camouflage according to the four combinations ways. The single-weather dataset in the table is identical to the multi-weather dataset except that the weather parameter is fixed to the default value in Carla. We observe that the multi-weather dataset and the NRP rendering component are crucial to improving attacks' performance. Specifically, the NRP rendering component has the most significant impact on adversarial camouflage effects, with an improvement of 0.34% under the multi-weather dataset.
We also observe that when the framework employs NR as the renderer, introducing multi-weather conditions for texture generation results in a slightly diminished attack performance. This discrepancy may stem from NR’s limitations in rendering comprehensive environmental characteristics. The weather information in the multi-weather dataset can’t be effectively incorporated into the foreground but instead amplifies the contrast between the foreground and background. Consequently, the image obtained from the fusion of the front and rear backgrounds becomes less realistic, causing the texture optimization in the wrong direction.
Table 5: The effectiveness of the multi-weather dataset for texture generation and NRP. Values are car AP@0.5 (%) with YOLOv3.

### v2

@@ 1 Introduction
Figure 1: Comparison of different adversarial camouflage under sunny (first row) and foggy (second row) environments, where only our method succeeds in both cases. (a) A car with normal texture; (b) DAS (Wang et al., 2021). (c) and (d) are top-performed methods FCA (Wang et al., 2022) and ACTIVE (Suryanto et al., 2023), respectively. (e) Our method RAUCA.
@@
Deep Neural Networks (DNNs) have achieved remarkable performance in many real-world applications such as face recognition and autonomous vehicles (Krizhevsky et al., 2012; Wang et al., 2023; Chai et al., 2021). However, DNNs suffer from adversarial examples (Szegedy et al., 2014). For instance, in the task of vehicle detection, adversarial inputs can deceive detection models, leading to incorrect detection of the surrounding vehicles, which poses a severe threat to the safety of autonomous vehicles.
@@
Adversarial examples can be categorized into digital and physical adversarial types. Digital adversarial examples involve introducing pixel-level perturbations to the model’s digital input, while physical ones manipulate real-world objects or surroundings, indirectly influencing the model inputs (Chen et al., 2018; Sharif et al., 2016a; Mahendran & Vedaldi, 2015). Physical adversarial examples are generally deemed more practical, as gaining direct access to the model’s input often necessitates system authentication. However, they are inherently more challenging as they must prove effective in complex physical environments, including various viewing perspectives, spatial distances, and lighting/weather conditions.
@@
This paper focuses on physical adversarial examples against vehicle (e.g., car) detection models, given their wide adoption in autonomous driving scenarios where safety is of great importance. To ensure the attack effectiveness across various viewing angles, current methods prefer generating adversarial camouflage capable of covering the entire surface of the vehicle (Wang et al., 2021, 2022; Suryanto et al., 2022, 2023). Top-performing methods achieve this by leveraging a differential neural renderer. This renderer maps the 3D vehicle and its texture to 2D images, establishing a differentiable path between the 3D vehicle and the vehicle detection models. Thus, the texture can be optimized through gradient back-propagation for effective camouflage generation.
@@
In the literature, there are two ways to generate camouflage with a neural renderer: one way is to optimize a 2D square texture pattern and project it onto the vehicle repeatedly, referred to as world-align-based methods; the other is to optimize the 3D texture of the vehicle in the form of UV-maps, referred to as UV-map-based method. However, both approaches currently suffer from certain issues. The former method cannot guarantee the texture pattern projected onto the car in the same manner in both texture generation and evaluation, leading to differences in adversarial camouflage between generation and evaluation. The latter UV-map-based render methods are not capable of rendering sophisticated environment characteristics such as light and weather, leading to sub-optimal camouflage generation. Besides, all current methods fail to consider the effectiveness of the camouflage under various weather conditions. As shown in Figure 1, the camouflage generated by the top-performed FCA (Wang et al., 2022) and ACTIVE (Suryanto et al., 2023) methods fails to attack the detector in a foggy scene.
@@
Our extensive studies demonstrate that our method outperforms the current state-of-the-art method by around 10% in car detection performance (e.g., AP@0.5). Our attacks achieve strong robustness and effectiveness, excelling in multi-weather, multi-view, and multi-distance conditions. We also demonstrate our method is effective in various weather conditions in the physical world. Our code is available at: https://github.com/SeRAlab/Robust-and-Accurate-UV-map-based-Camouflage-Attack.
@@ 2 Related work
Adversarial Camouflage: For self-driving cars, precise detection of surrounding vehicles is a critical safety requirement. Consequently, there has been a growing interest in developing adversarial vehicle camouflage to evade vehicle detection systems. Current research works mostly leverage a 3D simulation environment to obtain 2D rendered vehicle images with various transformations such that they can obtain robust adversarial camouflage. The early researches of adversarial camouflage for vehicles are mostly black-box methods because the rendering process in the simulation environment is non-differentiable. Zhang et al. (2019) first conducts experiments in the 3D space to generate adversarial camouflage for cars. They propose CAMOU, a method to train an approximate gradient network to mimic the behavior of both rendering and detection of the camouflage vehicles. Then, they can optimize the adversarial texture using this network. Meanwhile, Wu et al. (2020) propose an adversarial camouflage generation framework based on a genetic algorithm to search an adversarial texture pattern. Then, they repeat the optimized texture pattern to build the 3D texture that covers the full-body vehicle.
Recent methods introduce neural renderers, which enable differentiable rendering. With this technique, the adversarial texture can be optimized via gradient back-propagation. Two primary methods are employed in adversarial camouflage generation with neural renderer. One is to optimize a 3D model texture directly, which we refer to as UV-map-based camouflage methods; for example, Wang et al. (2021) propose a Dual Attention Suppression (DAS) attack, which minimizes the model attention and human attention on the camouflaged vehicle. Besides, Wang et al. (2022) propose the Full-coverage Camouflage Attack (FCA), which optimizes full-body surfaces of the vehicle in multi-view scenarios. The other is to optimize a 2D square texture pattern and then project it repeatedly to the target vehicle surface, which we refer to as world-align-based camouflage methods. Suryanto et al. (2022) present the Differentiable Transformer Attack (DTA), which proposes a differentiable renderer that can express physical and realistic characteristics (e.g., shadow). ACTIVE, proposed by Suryanto et al. (2023), introduces a texture mapping technique that utilizes depth images and improves the naturalness of the camouflage by using larger texture resolution and background colors.
@@ 3.1 Overview
Figure 2 shows our entire framework for adversarial camouflage generation. Firstly, we modify the weather parameters in the simulation environment to obtain a multi-weather vehicle dataset ($I_{in}$, $Y$, $\Phi_{cam}$, $M$), where $I_{in}$ is the original input images, $Y$ is the ground truth labels, $\Phi_{cam}$ is the camera pose parameters (position and angle) for viewing the car, and $M$ is the binary masks of $I_{in}$, where the target vehicle areas are set to 0. With $I_{in}$ and $M$, we can use:
$X_{ref}=I_{in}\cdot\left(1-{M}\right)$ (1)
$B=I_{in}\cdot{M}$ (2)
@@ 3.3 Neural Render Plus (NRP)
Before applying the framework to generate adversarial textures, NRP needs to be trained in advance to optimize the parameters of EFE. The NRP training process inputs the masked images of the white car $X_{ref}$, 3D mesh $Msh$, poses of the camera $\Phi_{cam}$, and multiple preset 3D texture T with different colors. At the same time, we need to obtain images of different vehicle colors based on $\Phi_{cam}$ and the colors of T from the photo-realistic renderer. Then, we cut out the vehicle parts as the network output target $TG$.
@@ 3.4 Attack Loss
where $x$ is the input image of the target detector, $H_{b}(x)$ is the detection bounding box, $gt$ is the ground-truth box and $IoU\left(H_{b}(x),gt\right)$ is the Intersection over Union (IoU) between $H_{b}(x)$ and $gt$. We use $IoU\left(H_{b}(x),gt\right)$ as a weight term, which allows the optimized loss function to focus more on the bounding box with a larger intersecting ratio area with $gt$. $H_{o}(x)$ and $H_{c}(x)$ represent the objectiveness score and the car class confidence score for the bounding box, respectively. $H_{d}\left.(x\right.)$ is our detection score, which is the product of the objectiveness confidence, class confidence, and the intersect ratio. Due to $IoU\left(H_{b}(x),gt\right)$, we only assign a non-zero detection score to the detected boxes which are intersected with the ground-truth box. This makes the texture optimization focus on making the target vehicle detection ineffective. We select the highest $H_{d}(x)$ and use it to compute $L_{atk}(x)$ through a log loss. By minimizing $L_{atk}(x)$, we can make the camouflaged vehicle misclassified or undetected by the object detector.
@@ 4.1 Experimental Settings
Datasets: We utilize CARLA to generate datasets in our experimentation. To facilitate a comparative analysis with previous studies (Wang et al., 2021, 2022; Suryanto et al., 2022, 2023), we select the Audi E-Tron as the target vehicle model. We create datasets using various simulation settings. For NRP training and testing, we utilized 69,120 and 59,152 photo-realistic images encompassing 16 distinct weather conditions, respectively. These conditions are a combination of four sun altitudes and four fog densities. Additionally, we generate a set of 40,960 images for our texture generation and prepare 7593 images for each adversarial camouflage in robust evaluation. The weather conditions present in these two datasets are the same conditions employed during the training and testing of NRP. For transferability evaluation, we utilize an “unseen-weather dataset” consisting of 7593 images in 16 novel weather conditions that are not seen during the texture generation phase. Moreover, we conduct experiments in the real world by printing five types of adversarial camouflages and sticking them to the 1:12 Audi E-Tron car models. We shoot 432 photos under different light conditions and 120 photos in a foggy environment for each model. Additional information regarding the construction process of the datasets can be found in the appendix.
@@ 4.2 Evaluation in Physically-Based Simulation Settings
Table 1: Comparison of the effectiveness of camouflages across various object detection models. Values are AP@0.5 (%) of the car.
@@
Figure 4: Attack comparison at night, a weather condition that has been included in our training set.
@@
Robustness in multi-weather dataset: We run an extensive attack comparison using diverse detection models. In addition to YOLOv3, we also use various black-box detection models to evaluate the camouflage. The results are shown in Table 1, showing that RAUCA has the best performance on all models. DAS generally performs less effectively, primarily due to the limitations of partially painted camouflage. Meanwhile, FCA exhibits sub-optimal performance, only slightly better than random camouflage, because it is not capable of rendering sophisticated environment characteristics during texture generation. DTA and ACTIVE are also less effective than our method because they do not consider multi-weather conditions during texture generation. Besides, their texture projection is not consistent between generation and testing. Our method demonstrates a nearly 13% improvement in attack effectiveness compared to other models on the white-box YOLOv3 model. Additionally, it achieves an improvement of around 8% on most other black-box models. Figure 5 shows an example of different vehicle camouflage at night: our method is still effective, while other methods fail to attack.
Figure 6 shows the summarized performance of each camera pose and weather parameter; values are car AP@0.5 averaged from the detectors used in Table 1. We can see that the texture produced by our method outperforms other approaches in most camera transformations and weather conditions. Our method improves multi-view robustness over previous methods in most viewpoints, thanks to the UV mapping-based projection for our textures and effective rendering of the environmental characteristics in the foreground. In addition, due to the incorporation of a multi-weather dataset during camouflage generation, our method is also the most robust in different weather environments.
@@
Transferability in unseen-weather dataset: We further evaluate the transferability of our camouflage under unseen weather conditions. As shown in Table 2, our method achieves the best attack performance on all types of detectors. Our method outperforms the previous state-of-the-art method, ACTIVE, by 23% on the white-box model YOLOv3, and approximately 11% on other black-box models. Our camouflage shows better attack performance under new weather conditions than under the weather conditions used for texture generation. This highlights the transferability of our approach across diverse weather conditions. Such good transferability stems from the use of the multi-weather dataset with various fog levels and lighting conditions for texture optimization. The dataset includes 16 different weather conditions, including extreme conditions, such as dark light and dense fog. In the dense fog weather, the vehicle camouflage is heavily obscured but still effective against detectors. The weather in real life is typically not so extreme, so our camouflage can transfer well and achieve effective attacking performance in the unseen conditions.
Figure 5 shows an example of different camouflage on rainy days. Notice that rainy weather is excluded from texture generation. Even though the rain blurs part of the texture, RAUCA uniquely succeeds in attacking the target detector.
@@ 4.3 Evaluation in Real-World Settings
We average the detection results for all environmental cases, and the results are in the last column of Table 3, which shows that our method has the best attack results. It illustrates the robustness of our method in different physical environments.
@@ 4.4 Ablation Studies
Effectiveness of the multi-weather dataset and NRP for texture generation: We evaluate the effectiveness of incorporating multi-weather conditions into our dataset and our proposed rendering component, NRP, for texture generation. As can be seen in Table 5, we generate the camouflage according to the four combinations ways. The single-weather dataset in the table is identical to the multi-weather dataset, except that the weather parameter is fixed to the default value in Carla. We observe that the multi-weather dataset and the NRP rendering component are crucial to improving attacks' performance. Specifically, the NRP rendering component has the most significant impact on adversarial camouflage effects, with an improvement of 0.34% under the multi-weather dataset.
We also observe that when the framework employs NR as the renderer, introducing multi-weather conditions for texture generation results in a slightly diminished attack performance. This discrepancy may stem from NR’s limitations in rendering comprehensive environmental characteristics. The weather information in the multi-weather dataset can’t be effectively incorporated into the foreground; instead, it amplifies the contrast between the foreground and background. Consequently, the image obtained from the fusion of the front and rear backgrounds becomes less realistic, causing the texture optimization in the wrong direction.
Effectiveness of our adversarial loss for texture generation: In this section, we evaluate the effectiveness of our designed adversarial loss function. We compare our adversarial loss with the ones used in FCA and ACTIVE. All these losses are calculated using the predicted detection boxes: the object confidence, class confidence, and IOU between the detection box and the ground truth. The difference between our designed loss function and FCA’s loss function is that FCA’s adversarial loss only considers the output detection box where the ground-truth center is located. In contrast, ours considers all of the detection boxes. It makes the camouflage to attack a broader range of detection boxes, preventing the target from being detected by boxes that do not contain the center of the ground truth. Compared with ACTIVE’s, our loss function incorporates the IOU value into the calculation, which makes our camouflage optimization focus more on confusing the detection boxes with a large degree of intersection with the target. Moreover, unlike minimizing the maximum confidence score across all classes in ACTIVE’s adversarial loss function, we minimize the maximum confidence score for the car class to achieve a more substantial attack effect. The comparing results are shown in Table 6. Values are the car AP@0.5. It can be seen that our proposed adversarial loss function achieves the strongest attack effect on most of the models.
Effectiveness of the smooth loss for texture generation: We investigate the impact of the smooth loss on the attack effectiveness of the generated camouflage. We vary the hyperparameter of the smooth loss $\beta$; the results are shown in Table 7. Values are the car AP@0.5. It can be seen that the overall attack effectiveness is better than the previous state-of-the-art ACTIVE method when $\beta$ changes, demonstrating the effectiveness of our generated camouflage.
Table 5: The effectiveness of the multi-weather dataset and NRP for texture generation. Values are car AP@0.5 (%) with YOLOv3.
@@
Table 6: The effectiveness of the different adversarial losses for texture generation. Values are car AP@0.5 (%) with YOLOv3.
Methods Single-stage Two-stage
YOLOv3 YOLOX DDTR DRCN SRCN FrRCN
FCA Loss 0.331 0.620 0.404 $\bm{0.437}$ 0.479 0.453
ACTIVE Loss 0.348 0.625 0.344 0.439 0.462 0.424
RAUCA Loss $\bm{0.304}$ $\bm{0.611}$ $\bm{0.285}$ 0.449 $\bm{0.384}$ $\bm{0.406}$
Table 7: The effectiveness of the smooth loss for texture generation. Values are car AP@0.5 (%) with YOLOv3.
Methods Single-stage Two-stage
YOLOv3 YOLOX DDTR DRCN SRCN FrRCN
ACTIVE 0.439 0.625 0.384 0.513 0.464 0.496
$\beta$=0.0 0.343 0.619 0.307 0.454 0.419 0.431
$\beta$=0.0001 $\bm{0.304}$ $\bm{0.611}$ $\bm{0.285}$ $\bm{0.449}$ $\bm{0.384}$ $\bm{0.406}$
$\beta$=0.0003 0.341 0.626 0.337 0.460 0.404 0.427